you inherit $16,000 and invest it in two bonds that pay 6% and 8% annual interest. At the end of one year the interest from the 8% bond is $580 greater than the interest from the 6% bond. How much was invested in each bond?
Answer by mananth(16949)

6% bond investment be x
8% bond investment be y
x+y = 16000.................(1)
6%x+580=8%y
multiply by 100
6x+58000=8y
6x-8y=-58000................(2)
solve simultaneously
1 x + 1 y = 16000 .............1

6 x -8 y = -58000 .............2
Eliminate y
multiply (1)by 8
Multiply (2) by 1
8 x 8 y = 128000
6 x -8 y = -58000
Add the two equations
14 x = 70000
/ 14
x = 5000
plug value of x in (1)
1 x + 1 y = 16000
5000 + y = 16000
y = 16000 -5000
y = 11000
y = 11000
x= 5000
y= 11000
$ 5000 @ 6%
$ 11000 @8%
m.ananth@hotmail.ca